OpenStudy (anonymous):

-5/6, -4/7, -3/8, -2/5 least to greatest

OpenStudy (cwrw238):

one way to do this would be to convert them to decimals eg -2/5 = -0.4 -5/6 = -0.833

OpenStudy (cwrw238):

it will then be obvious to arrange them least to greatest